Fatima is twice as old as her sister. In 5 years' time, she will be 2 years older than her sister. Make an equation and solve it to work out how old they both are now.

Answer by FrankM(1040)

F=2S
F+5=(S+5)+2
Substitute 2S for F in second equation
2S+5=(S+5)+2
2S+5=S+7
S=2
F=4